# An Exploratory Study on Using Large Language Models for Mutation Testing

### v1

@@ Abstract.
The question of how to generate high-utility mutations, to be used for testing purposes, forms a key challenge in mutation testing literature. Large Language Models (LLMs) have shown great potential in code-related tasks but their utility in mutation testing remains unexplored. To this end, we systematically investigate the performance of LLMs in generating effective mutations w.r.t. to their usability, fault detection potential, and relationship with real bugs. In particular, we perform a large-scale empirical study involving 4 LLMs, including both open- and closed-source models, and 440 real bugs on two Java benchmarks. We find that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs, which leads to approximately 18% higher fault detection than current approaches (i.e., 87% vs. 69%) in a newly collected set of bugs, purposely selected for evaluating learning-based approaches, i.e., mitigating potential data leakage concerns. Additionally, we explore alternative prompt engineering strategies and the root causes of uncompilable mutations, produced by the LLMs, and provide valuable insights for the use of LLMs in the context of mutation testing.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Mutation testing is one of the most effective software testing techniques (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). It is performed by producing program variants by making simple syntactic changes (i.e., mutations) in the source code of the origin program (i.e., mutants). These variants form the objectives of testing, in a sense requesting tests to trigger different behaviors on the original and the mutant programs. A mutant is called killed or live depending on whether the test execution leads to a different program output (observable behavior) from that of the original program or not. The proportion of the killed mutants over the entire set of mutants forms a test adequacy metric that is called mutation score.
@@
Unfortunately, rule-based approaches produce large numbers of redundant mutants, i.e., mutants that do not contribute to the testing process (Papadakis et al., 2016; Ammann et al., 2014) and thus exacerbate the method’s computational overheads as well as the human effort involved in examining live mutants. To address this issue, researchers leverage deep learning approaches to form mutation operators based on the project’s history (Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022), with the hope that these will produce few but effective mutations. While interesting, these approaches fail to produce effective mutations (Ojdanic et al., 2023b), perhaps due to the fact that they are trained on a few data instances.
@@
Nevertheless, the question of which mutants should be used when performing mutation testing remains open due to the following challenges: (1) Forming effective mutations. To be effective, mutations need to be syntactically correct, which is not the case of most of the existing learning-based methods, and be likely fault revealing (Chekam et al., 2020; Wen et al., 2019), i.e., have a good potential to couple with real faults. While existing methods are effective, there is still significant room for improvement (Degiovanni and Papadakis, 2022). (2) Forming natural mutations. Mutations need to align with natural coding patterns and practices so that they align with developer expectations, i.e., making them feel that they are worth being addressed (Beller et al., 2021). It is a challenge to ensure that the mutated code is not only syntactically correct but also aligns well with the realistic, common coding patterns and practices observed among developers. (3) Forming diverse and killable mutations. Determining whether a mutation is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021). At the same time, mutations need to be diverse and to cover the entire spectrum of code and program behaviours.
@@
To deal with these challenges, we aim at using Large-Language Models (LLMs), which have been trained on big code and can produce human-like code (Chen et al., 2021). Following existing studies (Guo et al., 2024; Ma et al., 2024; Schäfer et al., 2023), we designed a default prompt that contains task instructions, the code element to be mutated, the corresponding method code, and a few-shot examples sampled from real-world bugs. We carried out a large-scale evaluation involving 4 LLMs, 2 closed-source commercial and 2 open-source ones, 440 real-world bugs from two Java benchmarks, i.e., Defects4J (Just et al., 2014a) and ConDefects (Wu et al., 2023b), and compared with existing techniques and tools.
@@
Our evaluation includes both quantitative and qualitative analysis for the mutations generated by LLMs, comparing them with existing approaches to assess their cost and effectiveness. Additionally, we study how different prompt engineering strategies and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4-Turbo, CodeLlama-13b-Instruct (Roziere et al., 2023), and StarChat-$\beta$-16b (Li et al., 2023a)) affect the effectiveness of the task. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs.
Our results reveal that the best model for mutant generation is GPT-4 as it outperforms the other LLMs in all metrics used, i.e., produces less equivalent mutants, mutants with higher fault detection potential, as well as higher coupling and semantic similarity with real faults. GPT-3.5 is the second best in terms of fault detection potential, leading to a detection of 96.7% of bugs in Defects4J and 86.7% in ConDefects. From the rest of the techniques, Major is the best, with the potential to reveal 91.6% of bugs in Defects4J and 68.9% in ConDefects, indicating an advantage of approximately 18% for the GPT on ConDefects which is a newly and unseen dataset.
@@
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that GPTs exhibit the greatest diversity, newly introducing 45 different AST node types, in contrast to just 2 used by traditional approaches such as Major, and subsuming all existing approaches, indicating a much better contextualization of the LLM-based approach.
@@
We also analyzed the mutations generated by GPT that fail to compile, a common issue with learning-based mutations, and found that they fall in 9 compilation error types with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. This finding indicates that future approaches could focus on directing the LLMs in choosing the right method invocations when mutating.
@@
We wrapped up our method into a mutation tool called Kumo, an LLM-based mutation generation tool for Java. Based on the experimental data, we have also built a dataset that contains high-quality LLM-generated mutants that can be used not only in mutation testing but also in other bug-seeding applications, such as fault localization and fault prediction. The implementation of Kumo and the experimental data are available for open science (kum, 2024).
@@
  • We investigate the applicability of LLMs in mutation testing. We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our findings indicate that GPT models excel in generating mutations that closely mimic real bugs.
@@
  • We analyze the error types of non-compilable mutations and figure out that method invocation and member assessment are more likely to lead LLMs to generate non-compilable mutations.
  • We build a high-quality dataset of Java mutations that includes manually analysed equivalent mutants.
@@ 2. Background and Related Work
Mutation Testing and Mutation Generation.  Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978) one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Gligoric et al., 2013; Linares-Vásquez et al., 2017; Wang et al., 2022; Qian et al., 2022; Kazerouni et al., 2021; Groce et al., 2021; Wu et al., 2023a). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014b; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
All these tasks require high-utility mutations. This is because it is required to compile and execute every mutation with tests, which is computationally expensive. Therefore, generating redundant or non-compilable mutations, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2018; Wang et al., 2017, 2021; Garg et al., 2022). Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018). Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011; Zhu et al., 2023; Jiang et al., 2018), the behavior of the mutations needs to couple with real faults/bugs to allow these techniques to be effective.
@@
Several recent studies have also explored the use of LLMs for mutation generation (Tip et al., 2024; DEB et al., 2024). However, they have not been thoroughly evaluated, as they only count the number of mutations they introduce and compute the mutation scores achieved by some test suites. In contrast, we thoroughly assess the utility of the mutations we are using (through both a quantitative and qualitative approach), contrast it with the key previous baseline methods, and address all experimental parameters and concerns related to the use of LLMs, i.e., data leakage, cost, and added value.
@@
Closed-source LLMs are proprietary models developed and maintained by companies or organizations, often offering advanced features and requiring a subscription or payment for access, while open-source LLMs are freely available models that allow users to view and modify them. For the models used in our study, GPT-3.5-Turbo and GPT-4-Turbo are closed-source LLMs released by OpenAI (Achiam et al., 2023), while CodeLlama and StarChat are open-source. An open-source LLM is usually fine-tuned against a general LLM for specified tasks. For example, CodeLlama (Roziere et al., 2023) is derived from Llama (Touvron et al., 2023) by fine-tuning with extensive code datasets, while StarChat is built on StarCoder (Li et al., 2023a) to enhance conversational abilities.
@@ 3.1. Overview and Research Questions
  • RQ3: How do different prompts affect the performance in generating mutations?
  • RQ4: How do different LLMs affect the performance in generating mutations?
@@ 3.2. Datasets
ConDefects 45 2023/09/01 - 2023/09/30
Total 440 2006/07/16 - 2023/09/30
@@
To this end, we employ the Defects4J v1.2.0 (Just et al., 2014a) and ConDefects (Wu et al., 2023b) to evaluate the mutation generation approaches, shown as Table 1. In total, we conduct the experiments on 440 bugs.
Defects4J is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023a; Just et al., 2014b; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains history bugs from 6 open-source projects of diverse domains, ensuring a broad representation of real-world bugs. In total, these 6 projects contain 395 real bugs. However, from Table 2 and Table 1, we observe the time spans of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2023b), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder (atc, 2024) programming contest. To prevent data leakage, we exclusively use bugs reported after the LLMs’ release date, specifically those identified on or after August 31, 2023, and in total we collect 45 Java programs from ConDefects.
@@ 3.3. Mutation Generation via LLMs
3.3.1. Models
We aim to comprehensively compare the existing LLMs in their ability to generate mutations from various perspectives. This includes evaluating the capabilities of models that are fine-tuned on different base models and contrasting the performance of commercial closed-source models with open-source alternatives. In our study, we select LLMs based on the following criteria. (1) The LLM must have coding capabilities, i.e., understanding, generating, and transforming code. (2) The LLM must support interaction in chat mode, allowing for conversational code transformation. (3) The LLM must offer accessible APIs for integration.
Finally, we select the following models: GPT-3.5-Turbo (Achiam et al., 2023), GPT-4-Turbo (Achiam et al., 2023), CodeLlama-13b-Instruct (Roziere et al., 2023), and StarChat-$\beta$-16b (Li et al., 2023a)). The details of these models are shown as Table 2. These LLMs cover open-source (i.e., StarChat and CodeLlama) and closed-source (i.e., GPT-3.5-Turbo and GPT-4-Turbo), and various kinds of base models.

### v2

# On the Use of Large Language Models in Mutation Testing

@@ Abstract.
Mutation testing is a foundation approach in the software testing field, based on automatically seeded small syntactic changes, known as mutations. The question of how to generate high-utility mutations, to be used for testing purposes, forms a key challenge in mutation testing literature. Large Language Models (LLMs) have shown great potential in code-related tasks but their utility in mutation testing remains unexplored. To this end, we systematically investigate the performance of LLMs in generating effective mutations w.r.t. to their usability, fault detection potential, and relationship with real bugs. In particular, we perform a large-scale empirical study involving six LLMs, including both state-of-the-art open- and closed-source models, and 851 real bugs on two Java benchmarks (i.e., 605 bugs from 12 projects of Defects4J 2.0 and 246 bugs of ConDefects). We find that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs, which leads to approximately 19% higher fault detection than current approaches (i.e., 93% vs. 74%). Nevertheless, the mutants generated by LLMs have worse compilability rate, useless mutation rate, and equivalent mutation rate than those generated by rule-based approaches. This paper also examines alternative prompt engineering strategies and identifies the root causes of uncompilable mutations, providing insights for researchers to further enhance the performance of LLMs in mutation testing.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: International Conference on Automated Software Engineering; 27 October - 1 November, 2024; Sacramento, CA, USAisbn: 978-1-4503-XXXX-X/XX/XXjournal: JACMjournalvolume: 1journalnumber: FSEarticle: 1publicationmonth: 6ccs: Software and its engineering Software testing and debugging
@@ 1. Introduction
Mutation testing is one of the most widely-studied software testing techniques (Hamlet, 1977; DeMillo et al., 1978, 1979; Jia and Harman, 2010; Papadakis et al., 2019). It is performed by producing program variants by making simple syntactic changes (i.e., mutations) in the source code of the origin program (i.e., mutants). These variants form the objectives of testing, in a sense requesting tests to trigger different behaviors on the original and the mutant programs. A mutant is called killed or live depending on whether the test execution leads to a different program output (observable behavior) from that of the original program or not. The proportion of the killed mutants over the entire set of mutants forms a test adequacy metric that is called mutation score.
@@
Unfortunately, rule-based approaches produce large numbers of redundant mutants, i.e., mutants that do not contribute to the testing process (Papadakis et al., 2016; Ammann et al., 2014) and thus exacerbate the method’s computational overheads as well as the human effort involved in examining live mutants. To address this issue, researchers leverage deep learning approaches to form mutation operators based on the project’s history (Tufano et al., 2019; Patra and Pradel, 2021; Tian et al., 2022; Chen et al., [n. d.]), with the hope that these will produce few but effective mutations. While interesting, these approaches still have limitations in producing effective mutations (Ojdanic et al., 2023b), perhaps due to the fact that they are trained on a limited number of data instances.
@@
To date, the question of how to generate powerful mutants remains open due to the following challenges: (1) Forming effective mutations. To be effective, mutations need to be syntactically correct, which is not the case of most of the existing learning-based methods, and be likely fault revealing (Chekam et al., 2020; Wen et al., 2019), i.e., have a good potential to couple with real faults. While existing methods are effective, there is still significant room for improvement (Degiovanni and Papadakis, 2022). (2) Forming natural mutations. Mutations need to align with natural coding patterns and practices so that they align with developer expectations, i.e., making them feel that they are worth being addressed (Beller et al., 2021). It is a challenge to ensure that the mutated code is not only syntactically correct but also aligns well with the realistic, common coding patterns and practices observed among developers. (3) Forming diverse and killable mutations. Determining whether a mutation is equivalent is undecidable (Budd and Angluin, 1982), making it challenging to avoid the generation of equivalent mutants, leading to poor scalability and waste of resources (Zhang et al., 2013; Papadakis et al., 2015; Zhang et al., 2016; Wang et al., 2021). At the same time, mutations need to be diverse and cover the entire spectrum of code and program behaviors.
@@
This paper explores whether Large-Language Models (LLMs), trained on big code and can produce human-like code (Chen et al., 2021), could deal with these challenges. Following existing studies (Guo et al., 2024; Ma et al., 2024; Schäfer et al., 2023), we designed a default prompt that contains task instructions, the code element to be mutated, the corresponding method code, and a few-shot examples sampled from real-world bugs. We carried out a large-scale evaluation involving six LLMs (three closed-source commercial and three open-source ones), 851 real-world bugs from two Java benchmarks (i.e., 605 bugs across 12 projects from Defects4J 2.0 (Just et al., 2014a) and 246 bugs from ConDefects (Wu et al., 2024)), and compared with existing mutation testing techniques and tools.
@@
Our evaluation includes both quantitative and qualitative analysis of the mutations generated by LLMs, comparing them with existing approaches in terms of cost, usability (i.e., compilability rate, useless mutation rate, and equivalent mutation rate, AST node diversity), and behavior similarity with real bugs (i.e., real bug detection rate, coupling rate, and Ochiai Coefficient). Additionally, we study how different prompt engineering strategies and LLMs (i.e., GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o, GPT-4o-mini, DeepSeek-Coder-V2-236b (Zhu et al., 2024), StarChat-$\beta$-16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023)) affect the effectiveness of the task. Moreover, we analyze the root causes and error types for the non-compilable mutants generated by LLMs. Overall, our experiments generate about 7.8 million mutations and execute the origin test suites 5.7 million times.
Our results reveal that compared to existing approaches, LLMs generate more diverse mutations that are behaviorally closer to real bugs. This leads to higher fault detection rate than traditional approaches. For example, the mutants generated by GPT-4o lead to a fault detection rate of 93.4%, while the mutants generated by LEAM, PIT, and Major are 71.7%, 51.3%, and 74.4%, respectively. Nevertheless, we also observe that rule-based approaches, PIT and Major, outperform other approaches in terms of compilability rate, useless mutation rate, and equivalent mutation rate. In particular, Major excels with a 98.33% compilability rate, 0% useless mutations, and 0.59% equivalent mutants, while GPT4o exhibits a compilability rate of 75.6%, a useless mutation rate of 7.8%, and an equivalent mutation rate of 1.2%, respectively.
@@
Another interesting finding of our analysis regards the diversity of the mutants (measured in terms of newly introduced AST node types). In particular, our results show that GPT-4o-Mini and DeepSeek exhibit the greatest diversity, newly introducing 57 different AST node types. In comparison, rule-based approaches like Major (Just et al., 2011) introduce only 2 new node types, while LEAM (Tian et al., 2022), a state-of-the-art small model-based approach, introduces 31. The greater diversity, which subsumes all existing methods, suggests that LLMs can significantly diversify the mutations.
@@
We also analyzed the mutations generated by GPT models that fail to compile, the most common issue with learning-based mutations, and found that they fall in 9 compilation error types with the Usage of Unknown Methods and Code Structural Destruction types being the most prevalent ones. This highlights the need for improved strategies to guide model-generated mutations toward more syntactically valid and semantically meaningful variations, which could enhance the effectiveness of mutation testing using AI models.
@@
We wrapped up our LLM-based method into a mutation tool called Wukong. Based on the experimental data, we have also built a dataset that contains high-quality LLM-generated mutants that can be used not only in mutation testing but also in other bug-seeding applications, such as fault localization and fault prediction. The implementation of Wukong and the experimental data are available for open science (our, 2024).
@@
  • We investigate the applicability of LLMs in mutation testing. We perform extensive and detailed comparative experiments to evaluate LLMs against existing tools/methods. Our findings indicate that GPT-4 models excel in generating mutations that closely mimic real bugs.
@@
  • We analyze the error types of non-compilable mutations and figure out that member assessment and method invocation are more likely to lead LLMs to generate non-compilable mutations.
  • We build a high-quality dataset of Java mutations comprehensively labeled with detailed quality and behavior metrics.
@@ 2. Background and Related Work
Mutation Testing and Mutation Generation.  Mutation testing is a systematic testing approach that aims to guide the creation of high-utility test suites. It works by introducing syntactic modifications (i.e., mutations) into the source code of the programs under test. By observing if test suites can detect these changes (Hamlet, 1977; DeMillo et al., 1978) one can identify weaknesses of their test suites, i.e., mutants that are not killed. Besides testing, mutation analysis is a tool that can be used to support multiple applications in the field of software engineering (Shi et al., 2019; Linares-Vásquez et al., 2017). Most commonly, mutations are employed as substitutes for real-world bugs (Just et al., 2014b; Andrews et al., 2005; Daran and Thévenod-Fosse, 1996; Namin and Kakarla, 2011), to guide fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015), test prioritization (Shin et al., 2019) and program repair (Xiao et al., 2023, 2024).
All these tasks require high-utility mutations. This is because it is required to compile and execute every mutation with tests, which is computationally expensive. Therefore, generating redundant or non-compilable mutations, as done by most of the existing learning-based approaches, increases the computational cost involved (Zhang et al., 2018; Wang et al., 2017, 2021; Garg et al., 2022). Additionally, when mutations are used as substitutes for real defects, the mutations must be as syntactically close to the real bugs as possible, making the mutations more natural (Hindle et al., 2016; Jimenez et al., 2018; Hariri et al., 2019). Moreover, when performing tasks such as fault localization (Moon et al., 2014; Papadakis and Le Traon, 2015) and program repair (Le Goues et al., 2011), the behavior of the mutations needs to couple with real faults/bugs to allow these techniques to be effective.
@@
Several recent studies have also explored the use of LLMs for mutation generation. Tip et al. concurrently worked on using LLMs for mutation generation for JavaScript, but their approach neither included a comprehensive comparison with existing techniques nor did they utilize only three open-source LLMs (Tip et al., 2024). Similarly, Endres et al. employed large models for mutation generation to aid deduction (Endres et al., 2024), while Deb et al. used GPT as one of the baselines in their approach (DEB et al., 2024). In contrast to their projects, our work systematically explores a variety of LLMs and prompt designs, offering a thorough comparison with state-of-the-art approaches.
@@
Closed-source LLMs are proprietary models developed and maintained by companies or organizations, often offering advanced features and requiring a subscription or payment for access, while open-source LLMs are freely available models that allow users to view and modify them. For the models used in our study, GPT-3.5-Turbo, GPT-4o, and GPT-4o-Mini are closed-source LLMs released by OpenAI (Achiam et al., 2023), while DeepSeek-Coder, StarChat, and CodeLlama are open-source. An open-source LLM is usually fine-tuned against a base model for specified tasks. DeepSeek-Coder (Zhu et al., 2024) is based on DeepSeek. CodeLlama (Roziere et al., 2023) is derived from Llama (Touvron et al., 2023) by fine-tuning with extensive code datasets. StarChat is built on StarCoder (Li et al., 2023a) to enhance conversational abilities.
@@ 3.1. Overview and Research Questions
  • RQ3: How do different LLMs affect the performance in generating mutations?
  • RQ4: How do different prompts affect the performance in generating mutations?
@@ 3.2. Datasets
Cli 39 2007/05/15 - 2018/02/26
Codec 18 2008/04/27 - 2017/03/26
Csv 16 2012/03/27 - 2018/05/18
Gson 18 2010/11/02 - 2017/09/21
JacksonCore 26 2013/08/28 - 2019/04/05
Jsoup 93 2011/07/02 - 2019/07/04
ConDefects (CD) 246 2024/03/01 - 2024/06/30
Total 851 2006/07/16 - 2024/06/30
@@
To this end, we employ the Defects4J 2.0 (Just et al., 2014a) and ConDefects (Wu et al., 2024) to evaluate the mutation generation approaches, shown as Table 1. In total, we conducted experiments on 851 bugs.
Defects4J 2.0 is a widely used benchmark in the field of mutation testing (Ojdanic et al., 2023a; Just et al., 2014b; Kaufman et al., 2022; Degiovanni and Papadakis, 2022; Tian et al., 2022; Kim et al., 2022), which contains history bugs from open-source projects of diverse domains, ensuring a broad representation of real-world bugs. We select the most widely used 12 projects from the Defects4J 2.0, which contains 605 bugs in total. However, from Table 2 and Table 1, we observe the time spans of the Defects4J bugs are earlier than the LLMs’ training time, which may introduce data leakage. Therefore, we supplement another dataset, ConDefects (Wu et al., 2024), designed to address the data leakage concerns. ConDefects consists of tasks from AtCoder [^1] https://atcoder.jp programming contest. To prevent data leakage, we use data from the four most recent months following the release of the LLMs (i.e., from March 1, 2024, to June 30, 2024), during which we collected 246 Java bugs.
@@ 3.3. Mutation Generation via LLMs
3.3.1. Studied Models
We aim to comprehensively compare the existing LLMs in their ability to generate mutations from various perspectives. This includes evaluating the capabilities of models that are fine-tuned on different base models and contrasting the performance of commercial closed-source models with open-source alternatives. In our study, we select a representative set of LLMs with code capabilities. Following recent studies on evaluating the code capabilities of LLMs (Wang et al., 2024; Guo et al., 2024; Yuan et al., 2024), we adopt the closed-source LLMs from GPT family, including GPT-3.5-Turbo (Achiam et al., 2023), GPT-4o (Achiam et al., 2023), and GPT-4o-Mini (Achiam et al., 2023). Moreover, we employ the popular open-sourced LLMs that are used by recent studies (Ni et al., [n. d.]; Xia et al., 2024; Ma et al., 2024; Yang et al., 2024), including DeepSeek-Coder-V2-236b (Zhu et al., 2024), StarChat-$\beta$-16b (Li et al., 2023a), and CodeLlama-Instruct-13b (Roziere et al., 2023). The details of these models are shown as Table 2. Note that OpenAI has not officially released the parameter size of their closed-source models. The parameter size of the open-sourced models we studied ranges from 13 billion to 236 billion.